# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v1

@@ Appendix E Pseudocode for training with DeNS
We provide the pseudocode for training with DeNS in Algorithm 1 and note that Line 5 can be parallelized. For denoising partially corrupted structures discussed in Section 3.2.3, we only add noise to a subset of atoms (Line 16 – 19) and predict the corresponding noise (Line 30).
@@
4:$\sigma$: standard deviation of Gaussian noise if multi-scale noise is not used
5:$\sigma_{\text{high}}$: upper bound on standard deviations of Gaussian noise if multi-scale noise is used
6:$\sigma_{\text{low}}$: lower bound on standard deviations of Gaussian noise if multi-scale noise is used
7:$r_{\text{DeNS}}$: corruption ratio
8:$\lambda_{E}$: energy coefficient
9:$\lambda_{F}$: force coefficient
10:GNN: graph neural network for predicting energy, forces and noise
11:while training do
12:     $\mathcal{L}_{\text{total}}=0$
@@
13:     Sample a batch of $B$ structures $\{\left(S_{\text{non-eq}}\right)^{j}\mid j\in\{1,...,B\}\}$ from the training set
14:     for $j=1$ to $B$ do $\triangleright$ This for loop can be parallelized
15:         Let $\left(S_{\text{non-eq}}\right)^{j}=\left\{(z_{i},\mathbf{p}_{i})\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
@@
16:         Sample $p$ from a uniform distribution ${\mathbf{U}}(0,1)$ to determine whether to optimize DeNS
17:         if $p<p_{\text{DeNS}}$ then $\triangleright$ Optimize DeNS based on Equation 6
18:              if multi-scale noise is used then
19:                  Sample $\sigma_{\text{sample}}$ from $\{\sigma_{k}\}_{k=1}^{T}$
20:              else
21:                  $\sigma_{\text{sample}}=\sigma$
22:              end if
23:              for $i=1$ to $|\left(S_{\text{non-eq}}\right)^{j}|$ do
24:                  $q_{i}\sim{\mathbf{U}}(0,1)$
25:                  if $q_{i}<r_{\text{DeNS}}$ then $\triangleright$ Add noise to and denoise the atom
26:                       $\bm{\epsilon}_{i}\sim\mathcal{N}(0,\sigma_{\text{sample}}I_{3})$
27:                       $\tilde{\mathbf{p}_{i}}=\mathbf{p}_{i}+\bm{\epsilon}_{i}$
@@
28:                       $m_{i}=1$triangleright Denoise the atom when calculating $\mathcal{L}_{\text{DeNS}}$
29:                  else
30:                       $\tilde{\mathbf{p}_{i}}=\mathbf{p}_{i}$
31:                       $m_{i}=0$
32:                  end if
@@
33:                  $\tilde{\mathbf{f}}_{i}^{\prime}=\mathbf{f}_{i}^{\prime}\cdot m_{i}$triangleright Encode the atomic force if the atom is corrupted
34:              end for
35:              Let $(\tilde{S}_{\text{non-eq}})^{j}=\left\{(z_{i},\tilde{\mathbf{p}_{i}})\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
36:              Let $\tilde{F}\left((S_{\text{non-eq}})^{j}\right)=\left\{\tilde{\mathbf{f}}_{i}^{\prime}\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
37:              $\hat{E},\hat{F},\hat{\bm{\epsilon}}\leftarrow\text{GNN}\left((\tilde{S}_{\text{non-eq}})^{j},\tilde{F}\left((S_{\text{non-eq}})^{j}\right)\right)$
@@
38:              $\mathcal{L}_{E}=\left|E^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{E}\right|$
@@
39:              $\mathcal{L}_{\text{DeNS}}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}m_{i}\cdot\left|\frac{\bm{\epsilon}_{i}}{\sigma_{\text{sample}}}-\hat{\bm{\epsilon}}_{i}\right|^{2}$triangleright Calculate $\mathcal{L}_{\text{DeNS}}$ based on Equation 4
@@
40:              $\mathcal{L}_{F}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}(1-m_{i})\cdot|\mathbf{f}_{i}^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{\mathbf{f}}_{i}|^{2}$triangleright Predict forces for uncorrupted atoms
41:              $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{total}}+\lambda_{E}\cdot\mathcal{L}_{E}+\lambda_{\text{DeNS}}\cdot\mathcal{L}_{\text{DeNS}}+\lambda_{F}\cdot\mathcal{L}_{F}$
42:         elsenormal-$\triangleright$ Optimize the original task based on Equation 1
43:              $\hat{E},\hat{F},\_\leftarrow\text{GNN}\left((S_{\text{non-eq}})^{j}\right)$
44:              $\mathcal{L}_{E}=\left|E^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{E}\right|$
45:              $\mathcal{L}_{F}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}|\mathbf{f}_{i}^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{\mathbf{f}}_{i}|^{2}$
46:              $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{total}}+\lambda_{E}\cdot\mathcal{L}_{E}+\lambda_{F}\cdot\mathcal{L}_{F}$
47:         end if
48:     end for
49:     $\mathcal{L}_{\text{total}}=\frac{\mathcal{L}_{\text{total}}}{B}$
50:     Optimize GNN based on $\mathcal{L}_{\text{total}}$
51:end while
@@ Appendix F Visualization of corrupted structures
We visualize how adding noise of different scales affect structures in OC20, OC22 and MD17 datasets in Figure 3, Figure 4 and Figure 5, respectively.

### v2

@@ Appendix E Pseudocode for Training with DeNS
We provide the pseudocode for training with DeNS in Algorithm 1 and note that Line 5 can be parallelized. For denoising partially corrupted structures discussed in Section 3.2.3, we only add noise to a random subset of atoms (Line 11 – 14) and predict the corresponding noise (Line 25).
@@
4:$\sigma$: standard deviation of Gaussian noise
5:$r_{\text{DeNS}}$: corruption ratio
6:$\lambda_{E}$: energy coefficient
7:$\lambda_{F}$: force coefficient
8:GNN: graph neural network for predicting energy, forces and noise
9:while training do
10:     $\mathcal{L}_{\text{total}}=0$
@@
11:     Sample a batch of $B$ structures $\{\left(S_{\text{non-eq}}\right)^{j}\mid j\in\{1,...,B\}\}$ from the training set
12:     for $j=1$ to $B$ do $\triangleright$ This for loop can be parallelized
13:         Let $\left(S_{\text{non-eq}}\right)^{j}=\left\{(z_{i},\mathbf{p}_{i})\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
@@
14:         Sample $p$ from a uniform distribution ${\mathbf{U}}(0,1)$ to determine whether to optimize DeNS
15:         if $p<p_{\text{DeNS}}$ then $\triangleright$ Optimize DeNS based on Equation 6
16:              for $i=1$ to $|\left(S_{\text{non-eq}}\right)^{j}|$ do
17:                  $q_{i}\sim{\mathbf{U}}(0,1)$
18:                  if $q_{i}<r_{\text{DeNS}}$ then $\triangleright$ Add noise to and denoise the atom
19:                       $\bm{\epsilon}_{i}\sim\mathcal{N}(0,\sigma I_{3})$
20:                       $\tilde{\mathbf{p}_{i}}=\mathbf{p}_{i}+\bm{\epsilon}_{i}$
@@
21:                       $m_{i}=1$triangleright Denoise the atom when calculating $\mathcal{L}_{\text{DeNS}}$
22:                  else
23:                       $\tilde{\mathbf{p}_{i}}=\mathbf{p}_{i}$
24:                       $m_{i}=0$
25:                  end if
@@
26:                  $\tilde{\mathbf{f}}_{i}^{\prime}=\mathbf{f}_{i}^{\prime}\cdot m_{i}$triangleright Encode the atomic force if the atom is corrupted
27:              end for
28:              Let $(\tilde{S}_{\text{non-eq}})^{j}=\left\{(z_{i},\tilde{\mathbf{p}_{i}})\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
29:              Let $\tilde{F}\left((S_{\text{non-eq}})^{j}\right)=\left\{\tilde{\mathbf{f}}_{i}^{\prime}\mid i\in\{1,...,|\left(S_{\text{non-eq}}\right)^{j}|\}\right\}$
30:              $\hat{E},\hat{F},\hat{\bm{\epsilon}}\leftarrow\text{GNN}\left((\tilde{S}_{\text{non-eq}})^{j},\tilde{F}\left((S_{\text{non-eq}})^{j}\right)\right)$
@@
31:              $\mathcal{L}_{E}=\left|E^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{E}\right|$triangleright Predict energy of the original structure
@@
32:              $\mathcal{L}_{\text{DeNS}}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}m_{i}\cdot\left|\frac{\bm{\epsilon}_{i}}{\sigma}-\hat{\bm{\epsilon}}_{i}\right|^{2}$triangleright Predict noise of corrupted atoms
@@
33:              $\mathcal{L}_{F}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}(1-m_{i})\cdot|\mathbf{f}_{i}^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{\mathbf{f}}_{i}|^{2}$triangleright Predict forces of uncorrupted atoms
34:              $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{total}}+\lambda_{E}\cdot\mathcal{L}_{E}+\lambda_{\text{DeNS}}\cdot\mathcal{L}_{\text{DeNS}}+\lambda_{F}\cdot\mathcal{L}_{F}$
35:         $\triangleright$ Optimize the original task based on Equation 1
36:              $\hat{E},\hat{F},\_\leftarrow\text{GNN}\left((S_{\text{non-eq}})^{j}\right)$
37:              $\mathcal{L}_{E}=\left|E^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{E}\right|$
38:              $\mathcal{L}_{F}=\frac{1}{|(S_{\text{non-eq}})^{j}|}\sum_{i=1}^{|(S_{\text{non-eq}})^{j}|}|\mathbf{f}_{i}^{\prime}\left((S_{\text{non-eq}})^{j}\right)-\hat{\mathbf{f}}_{i}|^{2}$
39:              $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{total}}+\lambda_{E}\cdot\mathcal{L}_{E}+\lambda_{F}\cdot\mathcal{L}_{F}$
40:         end if
41:     end for
42:     $\mathcal{L}_{\text{total}}=\frac{\mathcal{L}_{\text{total}}}{B}$
43:     Optimize GNN based on $\mathcal{L}_{\text{total}}$
44:end while
@@ Appendix F Visualization of Corrupted Structures
We visualize how adding noise of different scales affects structures in OC20, OC22 and MD17 datasets in Figure 3, Figure 4 and Figure 5, respectively.